1. Rubber bungs are made by removing the tops of cones. Starting with a cone of radius 10cm and height 16cm, a rubber bungs made by cutting a cone of radius 5 cm and height 8cm from the top. Find the volume and total surface area of the rubber bung.
2. A metal cylinder is melted down and made into spherical balls for a game. The cylinder is 15cm high and has radius 6cm. The balls each have radius 1cm. How many balls can be made?

Answer by AnlytcPhil(1806)

1. Rubber bungs are made by removing the tops of cones. Starting
with a cone of radius 10cm and height 16cm, a rubber bungs made 
by cutting a cone of radius 5 cm and height 8cm from the top. 
Find the volume and total surface area of the rubber bung.

They start with a rubber cone whose cross section is like this:


             /|\
            / | \
         S /  |  \
          /   |16 \
         /    |    \
        /_____|_____\
           10

This has a volume of pr³/3 = p(10)³/3 = 1000p/3

Its surface area is found by
                          
Lateral Surface Area = prS, where S = slant height

(Note: Lateral Surface Area does not include the circular base)

Lateral Surface Area = prS = p(10)S = 10pS 
     _______    ___________    _______    ___    _____     __ 
S = Ör² + h² = Ö(10)²+(15)² = Ö100+225 = Ö325 = Ö25·13 = 5Ö13 
                                   __        __
Lateral Surface Area = prS = 10p(5Ö13) = 50pÖ13

Base (circle) Surface Area = pr² = p(10)² = 100p

Total Surface Area = Lateral surface area + Base surface area  
                         __               __
Total Surface Area = 50pÖ13 + 100p = 50p(Ö13 + 2)

------------------------------------------------------

Then they cut away a rubber cone whose cross section is like this:

             /|\
          S / |8\
           /__|__\
            5
This has a volume of pr³/3 = p(5)³/3 = 125p/3

Its surface area is found by
                          
Lateral Surface Area = prS, where S = slant height

Lateral Surface Area = prS = p(5)S = 5pS 
     _______    _________    _____    __    
S = Ör² + h² = Ö(5)²+(8)² = Ö25+64 = Ö89  
                                 __       __
Lateral Surface Area = prS = 5p(Ö89) = 5pÖ89

Note: They cut away only the lateral surface area of this small cone.
(Its base circle must then be added to find the total surface 
area of the bung)   

-----------------------------------------

They end up with a rubber bung whose cross section is like this:

           _5_____
          /   |   \
         /    |    \
        /_____|_____\
          10

Volume of bung = Volume of large cone - Volume of small cone  

Volume of bung = 1000p/3 - 125p/3 = 875p/3

That's approximately 916.3 cubic centimeters.

Total Surface Area of bung =
   Total Surface Area of large Cone - 
        Lateral Surface Area only of small Cone +
             Area of base of small cone (circular top of bung)

Area of circular top of bung = pr² = p(5)² = 25p
                                  __           __
Total Surface Area of bung = 50p(Ö13 + 2) - 5pÖ89 + 25p 
                                    __         __             
Total Surface Area of bung = 5p[10(Ö13 + 2) - Ö89 + 5]
                                   __         __             
Total Surface Area of bung = 5p[10Ö13 + 20 - Ö89 + 5]
                                   __    __             
Total Surface Area of bung = 5p(10Ö13 - Ö89 + 25)

That's approximately 810.9 square centimeters.
                                 
-----------------------------------------

Edwin